How do you put y= -8x^2+ 3 into vertex form?
Found 2 solutions by stanbon, lwsshak3:
Answer by stanbon(75887)

put y= -8x^2+ 3 into vertex form?
----
y = -8(x^2) + 3
-------
y = -8(x-0)^2 + 3
---------------------------
vertex: (0,3)
====================
Cheers,
Stan H.

Answer by lwsshak3(11628)

How do you put y= -8x^2+ 3 into vertex form?
**
Vertex form of equation for a parabola: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of the vertex, A is a multiplier which affects the slope or steepness of the curve.
..
Given equation is already in vertex form with:
vertex at (0,3) and A=-8